OpenStudy (anonymous):

If $535 is invested at an interest rate of 6% per year and is compounded continuously, how much will the investment be worth in 10 years?

jimthompson5910 (jim_thompson5910):

A=Pe^(rt) A=535*e^(0.06*10) A=535*e^(0.6) A=535*1.82211880039051 A=974.833558208922 A=974.83 So the account will be worth $974.83 in ten years.

OpenStudy (anonymous):

Thank you!<3

jimthompson5910 (jim_thompson5910):

you're welcome